OpenStudy (itsbribro):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary.

OpenStudy (itsbribro):



OpenStudy (itsbribro):

Line 1: y=-2x+____ Line 2: y=-2x+____

OpenStudy (itsbribro):

what is the blank or b

OpenStudy (itsbribro):

@acxbox22

OpenStudy (acxbox22):

find the y intercept of both lines the second line is easier because it is in the graph already

OpenStudy (itsbribro):

im sorry im still confused

OpenStudy (anonymous):

Find the slope of each line first, can you do that on your own?

OpenStudy (itsbribro):

i found the slopes i already put them in the line answers.

OpenStudy (acxbox22):

okay so line 1: find the slope and tell me line 2: slope is already given >>> (0,5)

OpenStudy (acxbox22):

okay you already have slope so for line we can substitue a point in the equation to find the y intercept

OpenStudy (acxbox22):

y=-2x+b 5=-2(-2)+b 5=4+b b=1 (0,1) is the y intercept

OpenStudy (itsbribro):

so for the blank for line 1 it would be 1

OpenStudy (itsbribro):

and for the blank for line 2 it would be 5?

OpenStudy (acxbox22):

yes